OpenStudy (anonymous):

A data set consists of the following data points: (2, 4), (4, 7), (5, 12) The slope of the best-fit line is 2.5. Find the y-intercept of this line. A. -1 B. -2.5 C. -1.5 D. -2

Mehek (mehek14):

let's use the first ordered pair \(4=2.5*2+b\) b is the y-intercept \(2.5*2=5\) \(4=5+b\) subtract 4-5 \(-1=b\)